OpenStudy (gabylovesyou):

@mathslover The formula for degrees Celsius is: C = 5/9 (F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is.

mathslover (mathslover):

Part 1: \[\large{C=\frac{5}{9}(F-32)}\] \[\large{C*9=\frac{5}{9}*9*(F-32)}\] can you tell me what will be the next step ?@Gabylovesyou

OpenStudy (gabylovesyou):

.....mmmmmm

mathslover (mathslover):

any confusion?

OpenStudy (gabylovesyou):

am i finding the solution for the F variable??

OpenStudy (gabylovesyou):

mmmmmm can u just explain it step by step cuz im really confused lol

mathslover (mathslover):

ok .. \[\large{9C=\frac{5}{\cancel{9}^1}*\cancel{9}^1*(F-32)}\]

mathslover (mathslover):

now divide both sides by 5: \[\large{\frac{9C}{5}=\frac{5(F-32)}{5}}\] \[\large{\frac{9C}{5}=\frac{\cancel{5}^1*(F-32)}{\cancel{5}^1}}\] \[\large{\frac{9C}{5}=F-32}\] \[\large{\frac{9C}{5}+32=F}\] got it?

OpenStudy (gabylovesyou):

yea for some reason when i first did it i got F = 1/5(9c + 160)

mathslover (mathslover):

well u r also right: \[\large{F=\frac{1}{5}(9C+160)}\] \[\large{F=\frac{9C}{5}+32}\]

OpenStudy (gabylovesyou):

they both equal??

OpenStudy (gabylovesyou):

now.... Part 2 :)